The electric field on the x axis due to a point charge fixed at the origin is given by Earrowbolditalic = (b/x2) ihatbold, where b = 5.00 kV·m and x > 0.
(a) Find the magnitude and sign of the point charge. (in nC)
(b) Find the potential difference between the points on the x axis at x = 1.50 m and x = 2.10 m.
Answer #1
E = b/x^2 = kq/x^2
b = kq
q = b/k = 5.00*10^3/8.99*10^9 = 5.56*10^-7 C
x1 = 1.50 m, x2 = 2.10 m
V1 - V2 = kq/x1 - kq/x2 = b(1/x1 - 1/x2) = 952 V